OpenStudy (anonymous):

What will the exponent be when you simplify the following problem? p^2,p^3=p^?

hartnn (hartnn):

why there is a \(\huge ,\) there ?? its is multiplication or division ?

hartnn (hartnn):

is it \(\large p^2 .p^3 =... ?\)

OpenStudy (anonymous):

doesn't say

OpenStudy (anonymous):

no its p^2,p^3=p^?

OpenStudy (anonymous):

thats exactly what it look like

hartnn (hartnn):

urgghhh...ambiguous question i will assume it is multiplication \(\large x^nx^m=x^{m+n}\) so what about \(\large p^2p^3=... ?\)

hartnn (hartnn):

basically you just add the exponents

OpenStudy (anonymous):

so its 5

hartnn (hartnn):

yup, \(p^2p^3 =p^5\)

OpenStudy (anonymous):

what about b^8/b^3=b^?

OpenStudy (anonymous):

do you divide 8 by 3

hartnn (hartnn):

when you divide, you SUBTRACT the exponents. \(\huge \dfrac{x^m}{x^n}=x^{m-n}\)

hartnn (hartnn):

so, what about b^8/b^3 ?

OpenStudy (anonymous):

so after dividing you subtract?

OpenStudy (anonymous):

5

hartnn (hartnn):

you don't divide the exponents you just subtract

hartnn (hartnn):

yes, b^8/b^3 = b^5 is correct

OpenStudy (anonymous):

ah ok thanks

hartnn (hartnn):

welcome ^_^